Problem 1 (#5-15, pg. 234): Determine the reactions at the supports. SkN 5 EN 78 KN
Answer #1

SIN Sy=0 Ay + BY = 6+8 = 14 kW AY Ax = 5 kn EMA = 0 By x6-1884). (6x2) - (582) 20 By=9kn Ay = 14-9 Slank